OpenStudy (anonymous):

What is the equation of the line tangent to the curve at P. Point P (0,1) and the curve that it lies on is √((x^(2)+y^(2))=tan((π/4)(x+y)). How do I begin? I tried implicit differentiation but I had too difficult of a time with it. I need to find the slope for y-1=m(x-0), right?

zepdrix (zepdrix):

Hey veronica :) Yes you have the right idea! Implicit differentiation giving you some trouble? Hmm let's see....

zepdrix (zepdrix):

\[\Large\rm \sqrt{x^2+y^2}=\tan\left[\frac{\pi}{4}(x+y)\right]\]This is the bad boy that we have to deal with? Did I copy that correctly?

zepdrix (zepdrix):

Do you remember your \(\Large\rm \sqrt{x}\) and \(\Large\rm \tan x\) derivatives?

zepdrix (zepdrix):

Vron?? D: Where hast thou gone?

OpenStudy (anonymous):

Oh sorry. I believe it is (1/2)x^(-1/2 and sec^(2)x, right?